OpenStudy (anonymous):

Please i REALLY need help with this ONE question Evaluate the following expression using the values given. Find 5x − 3y − z if x = −2, y = 2, and z = −3

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

5(-2)-3(2)-(-3)=-10-6+3=-13

OpenStudy (anonymous):

5(-2)-3(2)-(-3)=-10-6+3=-13

OpenStudy (anonymous):

thanks so much!